OpenStudy (anonymous):

Triangle ABC has side lengths 8, 15, and 17. Do the sides form a Pythagorean triple?

OpenStudy (anonymous):

\[17^2=289\\ 8^2+15^2=289=17^2\] So the sides form a Pythagorean triple !